OpenStudy (anonymous):

27 is what percent of 80? A. 33% B. 29.6% C. 33.75% D. 30%

OpenStudy (anonymous):

ok to do this you set it up as division 27/80 = ?

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

you need to solve x/100 = 27/80

OpenStudy (anonymous):

or 27/80 and multiply by 100